# A non-parametric approach for estimating consumer valuation distributions using second price auctions

### v2

@@ 3.4 The Coordinate ascent algorithm for maximizing $Lik_{A}$
As explained above, the adapted objective function $Lik_{A}$ and the search space $\mathcal{F}_{\bf z}$ of CDFs with relevant jump discontinuities are artifacts of the non-parametric approach that we pursue. However, once the estimates of the valuation distribution at elements of $\bf z$} are obtained using Algorithm 3.1, a continuous estimate on the entire valuation distribution can be constructed via interpolation. In particular, we use the values of $\widehat{F}_{(MLE)}$ at $z_{i}$’s, $\widehat{F}_{(MLE)}(0)=0$, and linear interpolation to construct a continuous estimator of the population valuation distribution over the entire real line.
@@ 4 Simulation study
In this section we consider various choices of the true underlying valuation distribution $F$, e.g., uniform, piecewise uniform, pareto, gamma, and beta distributions, which are commonly used in marketing research. We then illustrate and compare the performance of the MLE $\widehat{F}_{MLE}$, the initial estimate $\widehat{F}_{init}$ and the Polya tree estimate $\widehat{F}_{PT}$ estimate from George and Hui (2012) with the corresponding true valuation distribution $F$. Note that the Bayesian methodology in George and Hui (2012) (which uses only the final selling prices in each auction) requires the knowledge of the the total number of consumers accessing the auction. This does not hold in our motivating application, but in order to compare these methods we will also assume such a knowledge in our synthetic data evaluations for $\widehat{F}_{PT}$ below.
@@ 4.1 Data generation
From each of the five true underlying $F$’s, we consider two settings, with $K=100$ and $K=1000$ independent auctions of identical copies of an item. Varying the sample size here sheds light on the relationship between the sample size and the precision of the MLE $\widehat{F}_{MLE}$, the initial estimate $\widehat{F}_{init}$ and the Polya tree estimate $\widehat{F}_{PT}$ in estimating the true valuation distribution $F$. For each auction, we took the auction window ($\tau$) to be 100 units, and the constant rate ($\lambda$) of arrival of bidders to be equal to $1$. We then simulated the inter-arrival times between bidders from an exponential distribution with rate parameter $\lambda=1$, and drew the bidders’ valuations from $F$, keeping track of the entire sequence of standing prices and the intermediate times between jumps in the standing price for all independent auctions involved. This sequence of standing prices throughout the course of the auction, and the intermediate times between standing price changes are then treated as the observed dataset that is subsequently used to compute the initial estimator and the MLE. For each choice of true $F$ and number of auctions $K$, $100$ replicated datasets are generated.
Since the data is generated by consistent with the modeling assumptions, one expects $\widehat{F}_{MLE}$, which utilizes all available information, to have a superior performance than the initial estimator $\widehat{F}_{init}$, which only uses the final selling price and first non-reserve standing price for each auction, and the Polya tree estimator $\widehat{F}_{PT}$ of George and Hui (2012), which only uses final selling prices. The goal of these simulations is to examine extensively how much improvement can be obtained from our proposed method by incorporating the additional information in a variety of settings.
@@ 4.2 Simulation Results
For each replicated dataset generated (as described in the previous subsection), we apply our non-parametric methodology to obtain the initial estimate $\widehat{F}_{init}$, the MLE $\widehat{F}_{MLE}$, and the Polya Tree estimator $\widehat{F}_{PT}$ in George and Hui (2012). The goal is to compare the accuracy of each of these estimators with respect to the respective true valuation distribution $F$.
@@
Figure 5: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Polya Tree estimator (green) and the true valuation distribution $F$ (taken to be piece-wise Uniform) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Polya Tree) are also provided for both estimators (dotted lines, matching colors).
@@
Figure 6: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Polya Tree estimator (green) and the true valuation distribution $F$ (taken to be Uniform(1,20)(1,20)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Polya Tree) are also provided for both estimators (dotted lines, matching colors).
@@
Figure 7: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Polya Tree estimator (green) and the true valuation distribution $F$ (taken to be Pareto(location = 3,dispersion = 100)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Polya Tree) are also provided for both estimators (dotted lines, matching colors).
@@
Figure 8: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Polya Tree estimator (green) and the true valuation distribution $F$ (taken to be Gamma(shape = 10, rate = 2)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Polya Tree) are also provided for both estimators (dotted lines, matching colors).
@@
Figure 9: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Polya Tree estimator (green) and the true valuation distribution $F$ (taken to be Beta(2,2)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Polya Tree) are also provided for both estimators (dotted lines, matching colors).
@@
The above plots based on single chosen replicates are illustrative, but need to be complemented with performance evaluation averaged over all the $100$ replicates in each of the $10$ simulation settings. In Table 2, for each simulation setting, we provide both the KS-distance and the Total variation distance (TV-distance) between the true valuation distribution $F$ and the three estimates - $\widehat{F}_{MLE}$, $\widehat{F}_{init}$, $\widehat{F}_{PT}$ - averaged over the $100$ respective replications. The results show that the MLE (based on the entire collection of standing prices) uniformly outperforms the initial estimator (based only on final selling prices and first non-reserve standing prices) and the Polya Tree estimator (based only on final selling prices) in all the simulation settings. This strongly suggests that if the additional assumptions of single bidding and constant arrival rate seem to largely hold, it is worth using the proposed methodology which incorporates the extra information available in the form of all standing prices within the auction period.
KS distance Total Variation
Distribution $K$ MLE Initial Polya Tree MLE Initial Polya Tree
Uniform 100 0.048 0.138 0.050 0.075 0.211 0.058
1000 0.015 0.051 0.050 0.033 0.089 0.060
Piecewise Uniform 100 0.048 0.104 0.239 0.091 0.151 0.486
1000 0.015 0.035 0.237 0.065 0.082 0.479
Pareto 100 0.048 0.114 0.280 0.057 0.174 0.281
1000 0.017 0.041 0.251 0.019 0.069 0.249
Gamma 100 0.045 0.134 0.290 0.067 0.228 0.317
1000 0.014 0.051 0.296 0.022 0.083 0.309
Beta 100 0.054 0.146 0.121 0.074 0.237 0.133
1000 0.018 0.057 0.123 0.025 0.090 0.129
@@
Table 2: Kolmogorov-Smirnoff (KS) distance and Total variation distance between each of the three estimators $\widehat{F}_{MLE}$, $\widehat{F}_{init}$, Polya Tree estimator(PT) and the true valuation distribution $F$, averaged over 100 replications within each of the 10 simulation settings.
@@
For uncertainty quantification, we provide the coverage rate for 90% HulC-based intervals for the proposed MLE and Init approaches, and also the coverage rate for 90% credible intervals for the Polya Tree (PT) approach of George and Hui (2012), under varied simulation settings. Note that for every approach, we compute the coverage rate as the proportion of replications for which the entire true distribution function is contained within the respective upper and lower confidence bands (with Bonferroni correction used to adjust for multiple testing). The results show that the proposed MLE approach provides significantly better coverage than the other two approaches. We also computed the band area (area between the lower and upper confidence bands) for each of the approaches, and the respective values (averaged over all replications) are provided in Table 3 as well. These values reveal that the Init approach yields significantly wider confidence bands than the MLE approach, yet suffers from lower coverage. On the other hand, the PT approach yields significantly narrower bands compared to the MLE approach, but has zero empirical coverage. So overall, the proposed MLE approach strikes a much better balance overall than the other two estimators.
@@
The HulC approach assumes median unbiasedness of the underlying estimators, but $\widehat{F}_{init}$ and $\widehat{F}_{MLE}$ are not median unbiased - adjusting for median bias in HulC can be computationally expensive and leads to wider confidence bands. Note that near the boundaries, where the value of the underlying true CDF is very close to zero or one, median unbiasedness is a tall ask for these non-parametric estimators. Initial studies did show that the estimated median bias for $\widehat{F}_{MLE}$ was very small across the simulation settings for points away from the boundary (points within 5th to 95th percentile range). The last three columns of Table 3 provide ‘truncated coverage’ for the three methods across all settings, i.e., coverage calculation is restricted only to points within the 5th to 95th percentile range of the true distribution $F$. The results show that the observed coverage for $\widehat{F}_{init}$ and $\widehat{F}_{MLE}$ moves closer to the nominal coverage of 90% in many settings when we restrict away from the boundary. The comparative performance of the three methods remains the same. These results suggest that is not much advantage, in terms of coverage, by adjusting for median bias of the proposed MLE estimator in HulC.
@@ 5 Empirical application
In this section we apply our method to estimate the true valuation distribution of an Xbox based on actual data obtained from second-price auctions on eBay. In Section 5.1, we provide an overview of the data, and discuss features and adjustments to ensure its suitability for the methodology developed in the paper. In Section 5.2, we apply our non-parametric methodology on the data set and present the findings, and perform additional performance analysis.

### v3

@@ 3.4 The Coordinate ascent algorithm for maximizing $Lik_{A}$
As explained above, the adapted objective function $Lik_{A}$ and the search space $\mathcal{F}_{\bf z}$ of CDFs with relevant jump discontinuities are artifacts of the semi-parametric approach that we pursue. However, once the estimates of the valuation distribution at elements of $\bf z$} are obtained using Algorithm 3.1, a continuous estimate on the entire valuation distribution can be constructed via interpolation. In particular, we use the values of $\widehat{F}_{(MLE)}$ at $z_{i}$’s, $\widehat{F}_{(MLE)}(0)=0$, and linear interpolation to construct a continuous estimator of the population valuation distribution over the entire real line.
@@ 4 Simulation study
In this section we consider various choices of the true underlying valuation distribution $F$, e.g., uniform, piecewise uniform, pareto, gamma, and beta distributions, which are commonly used in marketing research. We then illustrate and compare the performance of the MLE $\widehat{F}_{MLE}$, the initial estimate $\widehat{F}_{init}$ and the Pólya Tree estimate $\widehat{F}_{PT}$ estimate from George and Hui (2012) with the corresponding true valuation distribution $F$. Note that the Bayesian methodology in George and Hui (2012) (which uses only the final selling prices in each auction) requires the knowledge of the total number of consumers accessing the auction. This does not hold in our motivating application, but in order to compare these methods we will also assume such a knowledge in our synthetic data evaluations for $\widehat{F}_{PT}$ below.
@@ 4.1 Data generation
From each of the five true underlying $F$’s, we consider two settings, with $K=100$ and $K=1000$ independent auctions of identical copies of an item. Varying the sample size here sheds light on the relationship between the sample size and the precision of the MLE $\widehat{F}_{MLE}$, the initial estimate $\widehat{F}_{init}$ and the Pólya Tree estimate $\widehat{F}_{PT}$ in estimating the true valuation distribution $F$. For each auction, we took the auction window ($\tau$) to be 100 units, and the constant rate ($\lambda$) of arrival of bidders to be equal to $1$. We then simulated the inter-arrival times between bidders from an exponential distribution with rate parameter $\lambda=1$, and drew the bidders’ valuations from $F$, keeping track of the entire sequence of standing prices and the intermediate times between jumps in the standing price for all independent auctions involved. This sequence of standing prices throughout the course of the auction, and the intermediate times between standing price changes are then treated as the observed dataset that is subsequently used to compute the initial estimator and the MLE. For each choice of true $F$ and number of auctions $K$, $100$ replicated datasets are generated.
Since the data is generated to be consistent with the modeling assumptions, one expects $\widehat{F}_{MLE}$, which utilizes all available information, to have a superior performance than the initial estimator $\widehat{F}_{init}$, which only uses the final selling price and first non-starting standing price for each auction, and the Pólya Tree estimator $\widehat{F}_{PT}$ of George and Hui (2012), which only uses final selling prices. The goal of these simulations is to examine extensively how much improvement can be obtained from our proposed method by incorporating the additional information in a variety of settings.
@@ 4.2 Simulation Results
For each replicated dataset generated (as described in the previous subsection), we apply our semi-parametric methodology to obtain the initial estimate $\widehat{F}_{init}$, the MLE $\widehat{F}_{MLE}$, and the Pólya Tree estimator $\widehat{F}_{PT}$ in George and Hui (2012). The goal is to compare the accuracy of each of these estimators with respect to the respective true valuation distribution $F$.
@@
Figure 5: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Pólya Tree estimator (green) and the true valuation distribution $F$ (taken to be piece-wise Uniform) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Pólya Tree) are also provided for both estimators (dotted lines, matching colors).
@@
Figure 6: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Pólya Tree estimator (green) and the true valuation distribution $F$ (taken to be Uniform(1,20)(1,20)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Pólya Tree) are also provided for both estimators (dotted lines, matching colors).
@@
Figure 7: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Pólya Tree estimator (green) and the true valuation distribution $F$ (taken to be Pareto(location = 3,dispersion = 100)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Pólya Tree) are also provided for both estimators (dotted lines, matching colors).
@@
Figure 8: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Pólya Tree estimator (green) and the true valuation distribution $F$ (taken to be Gamma(shape = 10, rate = 2)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Pólya Tree) are also provided for both estimators (dotted lines, matching colors).
@@
Figure 9: Plot of the MLE $\widehat{F}_{MLE}$ (red), initial estimator $\widehat{F}_{init}$ (blue), Pólya Tree estimator (green) and the true valuation distribution $F$ (taken to be Beta(2,2)) for a random chosen replicate with $K=100$ (left) and $K=1000$ independent auctions (right). $90\%$-HulC confidence intervals or credible interval (for Pólya Tree) are also provided for both estimators (dotted lines, matching colors).
@@
The above plots based on single chosen replicates are illustrative, but need to be complemented with performance evaluation averaged over all the $100$ replicates in each of the $10$ simulation settings. In Table 2, for each simulation setting, we provide both the Kolmogorov-Smirnov distance (KS-distance) and the Total variation distance (TV-distance) between the true valuation distribution $F$ and various estimators. In this comparison, we also include two parametric methods, the Gamma-F method and the Truncated Normal-F (TN-F) method, used in George and Hui (2012). Both methods use final selling prices, and similar to Pólya Tree, require the knowledge of the total number of potential bidders who access the auction. The distances for all five methods (MLE, Initial, Pólya Tree, Gamma-F and TN-F) are averaged over the $100$ respective replications. The results show that the MLE (based on the entire collection of standing prices) uniformly outperforms the initial estimator (based only on final selling prices and first non-starting standing prices) and the other three final selling price based estimators (Pólya Tree, Gamma-F and TN-F) in all the simulation settings. This strongly suggests that if the additional assumptions of single bidding and constant arrival rate seem to largely hold, it is worth using the proposed methodology which incorporates the extra information available in the form of all standing prices within the auction period.
KS distance
Distribution $K$     MLE Initial Pólya Tree Gamma-F TN-F
Uniform 100 0.048 0.138 0.050 0.266 0.732
1000 0.015 0.051 0.050 0.302 0.727
Piecewise-Uniform 100 0.048 0.104 0.239 0.385 0.708
1000 0.015 0.035 0.237 0.397 0.716
Pareto 100 0.048 0.114 0.280 0.061 0.131
1000 0.017 0.041 0.251 0.021 0.124
Gamma 100 0.045 0.134 0.290 0.314 0.239
1000 0.014 0.051 0.296 0.348 0.229
Beta 100 0.054 0.146 0.121 0.281 0.488
1000 0.018 0.057 0.123 0.382 0.478
Total Variation
Uniform 100 0.075 0.211 0.058 0.296 0.738
1000 0.033 0.089 0.060 0.313 0.734
Piecewise-Uniform 100 0.091 0.151 0.486 0.445 0.674
1000 0.065 0.082 0.479 0.433 0.691
Pareto 100 0.057 0.174 0.281 0.057 0.13
1000 0.019 0.069 0.249 0.021 0.123
Gamma 100 0.067 0.228 0.317 0.316 0.241
1000 0.022 0.083 0.309 0.351 0.229
Beta 100 0.074 0.237 0.133 0.288 0.483
1000 0.025 0.090 0.129 0.385 0.472
@@
Table 2: Kolmogorov-Smirnov (KS) distance and Total variation distance between various estimators ($\widehat{F}_{MLE}$, $\widehat{F}_{init}$, Pólya Tree estimator(PT), the Gamma method based on final selling prices - Gamma-F, the Truncated-Normal method based on final selling prices - TN-F) and the true valuation distribution $F$, averaged over 100 replications within each of the 10 simulation settings.
@@
For uncertainty quantification, we provide the coverage rate for 90% HulC-based intervals for the proposed MLE and Init approaches, and also the coverage rate for 90% credible intervals for the Pólya Tree (PT) approach of George and Hui (2012), under varied simulation settings. Note that for every approach, we compute the coverage rate as the proportion of replications for which the entire true distribution function is contained within the respective upper and lower confidence bands (with Bonferroni correction used to adjust for multiple testing). The results show that the proposed MLE approach provides significantly better coverage than the other two approaches. We also computed the band area (area between the lower and upper confidence bands) for each of the approaches, and the respective values (averaged over all replications) are provided in Table 3 as well. These values reveal that the Init approach yields significantly wider confidence bands than the MLE approach, yet suffers from lower coverage. On the other hand, the PT approach yields significantly narrower bands compared to the MLE approach, but has zero empirical coverage. So overall, the proposed MLE approach strikes a much better balance overall than the other two estimators.
@@
The HulC approach assumes median unbiasedness of the underlying estimators, but $\widehat{F}_{init}$ and $\widehat{F}_{MLE}$ are not median unbiased - adjusting for median bias in HulC can be computationally expensive and leads to wider confidence bands. Note that near the boundaries, where the value of the underlying true CDF is very close to zero or one, median unbiasedness is a tall ask for these non-parametric/semi-parametric estimators. Initial studies did show that the estimated median bias for $\widehat{F}_{MLE}$ was very small across the simulation settings for points away from the boundary (points within 5th to 95th percentile range). The last three columns of Table 3 provide ‘truncated coverage’ for the three methods across all settings, i.e., coverage calculation is restricted only to points within the 5th to 95th percentile range of the true distribution $F$. The results show that the observed coverage for $\widehat{F}_{init}$ and $\widehat{F}_{MLE}$ moves closer to the nominal coverage of 90% in many settings when we restrict away from the boundary. The comparative performance of the three methods remains the same. These results suggest that is not much advantage, in terms of coverage, by adjusting for median bias of the proposed MLE estimator in HulC.
@@
4.3 Evaluation based on profit-maximizing price
As noted in George and Hui (2012), a key advantage of knowing the valuation distribution $F$ is that it enables managers to determine the profit-maximizing price for a product in a given market. In particular, if $c$ is the production cost, then the expected (per person) profit at a price $x$ is given by
$\pi(x)=(1-F(x))(x-c)$
In particular, $\pi(x)$ is the total expected profit divided by the total number of potential customers in the underlying population. As George and Hui (2012) note, $\pi(x)$ is the product of the proportion of potential customers with valuation above price $x$ (i.e., 1F(x)1-F(x)), and the profit for each sale (i.e., xcx-c). Hence, the profit maximizing price $x_{F}$, is defined as
$x_{F}=\mbox{argmax}_{x}(1-F(x))(x-c).$
Based on an estimate $\widehat{F}$ of $F$, the underlying profit maximizing price $x_{F}$ can be estimated by
$x_{\widehat{F}}=\mbox{argmax}_{x}(1-\widehat{F}(x))(x-c).$
Based on these considerations, for each of the ten simulation settings—corresponding to five choices of the true valuation distribution $F$ and two choices of the total number of auctions KK—we now additionally evaluate each method along two dimensions: (a) the proximity of the estimated profit‑maximizing price $x_{\widehat{F}}$ to the true profit‑maximizing price $x_{F}$, and (b) the proximity of the estimated (per person) maximum profit
$\widehat{\pi}(x_{\widehat{F}})=\bigl(1-\widehat{F}(x_{\widehat{F}})\bigr)\bigl(x_{\widehat{F}}-c\bigr)$
[…]
We use five methods in our comparison. These include the proposed MLE approach, the Initial approach, the Pólya Tree approach of George and Hui (2012) and two parametric approaches based on final selling prices used in George and Hui (2012) - the Gamma-F method, and the Truncated Normal-F (TN-F) method.
Figure 10 provides box plots for the estimated (per person) profit maximizing price $x_{\widehat{F}}$ of each method in each of the 10 simulation settings. The production cost $c$ was set to the 25% quantile of the underlying true distribution, and the results described below exhibit similar patterns even when $c$ varies from the 10% to the 50% quantile (see Supplemental Section S5 for more details).
It can be seen that the estimated profit‑maximizing price produced by the MLE method is generally very close to the true optimal price, while exhibiting lower variability than the Initial method. In contrast, the Pólya tree method tends to overestimate the optimal price across nearly all settings (Piecewise Uniform being the sole exception), whereas the Gamma method typically underestimates it (Gamma being the sole exception). The Truncated-Normal method tends to have a much larger bias compared to the other methods.
[…]
In contrast, the Gamma method systematically underestimates the maximum profit in nearly all settings, with the Pareto case being the sole exception. The Pólya tree estimator displays more variable behavior: it tends to overestimate profits most severely in the Pareto setting, while underestimating profits in the Uniform and the Piecewise Uniform settings. The Truncated-Normal method performs the worst overall, showing poor accuracy relative to the other estimators in most simulation settings.
Figure 10: Box plots for the estimated profit maximizing price $x_{\widehat{F}}$ for each of the five estimators, $\widehat{F}_{MLE}$, $\widehat{F}_{init}$, Pólya tree estimator (PT), the Gamma method based on final selling prices (Gamma-F), and the Truncated-Normal method based on final selling prices (TN-F). Results are based on 100 replications within each of the 10 simulation settings. The dashed lines represent the true profit maximizing price $x_{F}$ in each scenario.
[…]
Distribution $K$     MLE Initial Pólya Tree Gamma-F TN-F
Uniform 100 +0.064 +0.235 -0.119 -1.094 +7.250
1000 +0.022 +0.115 -0.096 -1.286 +7.164
Piecewise-Uniform 100 +0.006 +0.007 -0.240 -0.157 +1.399
1000 +0.001 +0.008 -0.242 -0.192 +1.426
Pareto 100 +0.027 +0.191 +0.915 +0.033 +0.370
1000 +0.009 +0.055 +0.739 -0.006 +0.347
Gamma 100 +0.013 +0.118 +0.022 -0.256 -0.117
1000 +0.003 +0.026 -0.014 -0.277 -0.117
Beta 100 +0.002 +0.017 -0.003 -0.050 -0.071
1000 +0.000 +0.004 -0.005 -0.063 -0.071
[…]
@@ 5 Empirical application
In this section we apply our method to estimate the true valuation distribution of an Xbox based on actual data obtained from second-price auctions on eBay. In Section 5.1, we provide an overview of the data, and discuss features and adjustments to ensure its suitability for the methodology developed in the paper. In Section 5.2, we apply our semi-parametric methodology on the data set and present the findings, and perform additional performance analysis.